OpenStudy (anonymous):

For the parent function y equals the square root of x, what effect does a value of b = -2 have on the graph? A. Horizontal compression by a factor of one half and a reflection across the y-axis B. Horizontal compression by a factor of one half and a reflection across the x-axis C. Horizontal compression by a factor of negative one half and a reflection across the y-axis D. Horizontal compression by a factor of negative one half and a reflection across the x-axis

OpenStudy (anonymous):

@ceecee_millions is it C?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so the equation is \[y = \sqrt{x}\] where does the b go...? 1. \[y = \sqrt{x} + b\] or 2. \[y = \sqrt{x + b}\]

OpenStudy (campbell_st):

but a really good method for this is to use some graphing software and look at things yourself. I'd recommend https://www.desmos.com/calculator

OpenStudy (anonymous):

2 right?

OpenStudy (anonymous):

@EmilyF03 what do you think?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

is it C @campbell_st